Circular No. 7213

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1999 N1 (SOHO) D. A. Biesecker, SM&A Corporation and Goddard Space Flight Center,

reports observations of another apparent Kreutz sungrazing comet discovered by Darren Lewis (University of Birmingham) in the SOHO/LASCO data. The comet is visible in both the C2 and C3 telescopes, and no tail is evident. Astrometric measurements made by Biesecker (and reduced by B. G. Marsden), together with orbital elements by Marsden, are given on MPEC 1999-N02.

1999 UT  α_{2000}  δ_{2000}
July 1.071  6^{h}30.^{m}6  +21^{o}10′

GRB 990627 L. Nicastro, Istituto di Fisica Cosmica ed Applicazioni all’Informatica,

CNR, Palermo; L. A. Antonelli, Osservatorio Astronomico, Rome; M. Dadina and M. R. Daniele, BeppoSAX Science Data Center, Rome; E. Costa, Istituto di Astrofisica Spaziale, CNR, Rome; and E. Pian, Istituto di Tecnologie e Studio Radiazioni Extraterrestri, CNR, Bologna, report: “The BeppoSAX Wide Field Camera (WFC) error box of GRB 990627 (IAUC 7211, GCN 356) was observed from June 27.54 (8 hr after the γ-ray trigger) to 28.86 UT, using the Narrow Field Instruments of BeppoSAX. Two weak, previously unknown sources designated as 1SAX J0148.5–7704 (α = 1^{h}48^{m}27^{s}, δ = −77^{o}04′.6, equinox 2000.0; error radius 1′) and 1SAX J0149.1–7709 (α = 1^{h}49^{m}13^{s}, δ = −77^{o}08′.9) were detected by the MECS units. 1SAX J0149.1–7709 is outside of the 3′ WFC error-circle (4′.5 from the WFC centroid), and its flux does not vary during the observation. 1SAX J0148.5–7704 lies at 43′′ from the WFC centroid and fades by a factor three or more. Its mean 2–10-keV flux during the first half of the observation is 3.5 × 10^{−13} erg cm^{−2} s^{−1} (assuming a Crab-like energy spectrum and Galactic value for the electron column density), and it is no longer detectable in the second half. The 1SAX J0148.5–7704 position is compatible with the radio source AT J014832.6–770412 reported by Subrahmanyan et al. (GCN 357). We conclude that this source is the x-ray afterglow of GRB 990627.”

1999 July 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green